OpenStudy (anonymous):

Can someone help me with this? 3x+4 over 3 plus 2x over 3-x = x

OpenStudy (aravindg):

can u pls use eqn editor?

OpenStudy (anonymous):

(3x+4)/(3+2x)/(3-x)= x Is this it?

OpenStudy (anonymous):

Wait wait

OpenStudy (anonymous):

multiply every term by 12, 8x-12 + 12 = 9x+15-24 x= 9 SOURCE: http://answers.yahoo.com/question/index?qid=20080915034250AA7J4yD

OpenStudy (aravindg):

did u mean \[\large \frac{3x+4}{3}+\frac{2x}{3-x}=x\;\;\;\;?\]

OpenStudy (anonymous):

I don't think yahoo answers is a reliable source

OpenStudy (anonymous):

\[\left\{ 3x+4 \over 3 \right\} + \left\{ 2x \over 3-x \right\} = x\]

OpenStudy (aravindg):

i agree @Romero

OpenStudy (aravindg):

:) i gt qn right

jhonyy9 (jhonyy9):

chol so first you need calcule the common denominator

OpenStudy (anonymous):

Try to get the denominator the same for both of them?

OpenStudy (anonymous):

How will I solve it?

OpenStudy (anonymous):

Yeah, but I don,t know how to cancel and multiply or sorts of that.

jhonyy9 (jhonyy9):

how you do in case of 1/2 + 3/5 = ?

OpenStudy (anonymous):

x=-6 work it and you will find it

OpenStudy (anonymous):

You sure with x= -6? I kind of check it, but it got wrong

OpenStudy (anonymous):

That's why you don't take answers from someone who doesn't show their work.

jhonyy9 (jhonyy9):

chol you dont like getting the denominator common ?

OpenStudy (anonymous):

cross multiply to get\[\frac{(3x+4)(3-x) + 3(2x)}{3(3-x)}\]

OpenStudy (anonymous):

Yes try getting the common denominator

OpenStudy (anonymous):

@psujono Try to let the asker solve the question.

OpenStudy (anonymous):

i am sure one billion percent

OpenStudy (anonymous):

@jhonyy9 Because I don't know how to cancel or multiply when it comes to rational expressions with the common denominator

OpenStudy (anonymous):

Sorry @Romero

OpenStudy (anonymous):

red123url stop giving answers if you don't know how to solve it.

OpenStudy (anonymous):

come on I know the answer but I want her to solve it herself and I am lazy I dont want to write equations here

OpenStudy (anonymous):

I also got -6 when I answered earlier but when I checked it, it was not equal

OpenStudy (anonymous):

That's very dumb. Why would anyone bother solving it once they are given an answer? And no one really cares if you know the answer.

OpenStudy (anonymous):

Try solving by getting a common denominator

OpenStudy (anonymous):

here we are trying to help people not to give them everything and by the way u r rude

OpenStudy (anonymous):

red what is this?

OpenStudy (anonymous):

are talking about me swalah ?

OpenStudy (anonymous):

-3 is not check either

OpenStudy (anonymous):

Wrong, wrong. Wait

OpenStudy (anonymous):

what is wrong with -6

OpenStudy (anonymous):

Try to check it with the equation, it is not equal

jhonyy9 (jhonyy9):

chol - sorry but now i remember that when you need assum fractions you need push first of all restrictions for x when for this value of x the denominator will be zero so because than the farction is undefined

jhonyy9 (jhonyy9):

sorry the fraction is undefined

OpenStudy (anonymous):

no it is, (3(-6)+4)/3+(2(-6))/3-(-6) -18+4/3+-12/9 -14/3+(-4)/3= -18/3=-6 which is x

jhonyy9 (jhonyy9):

hope so much that is understandably sure right

jhonyy9 (jhonyy9):

- so than x not can being equal ?

OpenStudy (anonymous):

Sorry @RED123url, it's correct

OpenStudy (anonymous):

some people want to report me I dont know why ??? maybe cus I give answers :(